# Enhancing LLM Reasoning with Iterative DPO: A Comprehensive Empirical Investigation

### v1

@@ 1 Introduction
    Iterative Improvement of Generator and RM can be Achieved via Multi-Round DPO: By relabeling the RM with online-generated data, the mutual enhancement of the generator and RM can be realized through multiple rounds of DPO — a capability that is often challenging to attain in traditional RL frameworks.
@@
    Multi-Round DPO Achieves RL-Level Performance with Lower Computational Cost: Our final model DPO-VP, based on Qwen2.5-Math-7B, achieves comparable average performance to SimpleRL-Zero and PURE-VR across five challenging mathematical benchmarks (48.2 vs 48.8/47.7), while requiring only a single 80GB GPU for training.
@@ 3.5.1 Setup and Main Results
To align with existing RL methods, following Zeng et al. (2025) and Cheng et al. (2025), we select 8K Level 3-5 questions from the MATH as self-improvement data, and additionally evaluate on Olympiad-level benchmarks, including AMC23, AIME24, and OlympiadBench He et al. (2024). Furthermore, we conduct additional comparisons on two out-of-domain datasets to assess generalization performance. More training details can be found in Appendix D and E. Table 2 presents the pass@1 accuracy of different models, while Table 3 further analyzes training data consumption and GPU usage.
@@
    Multi-round DPO can achieve mathematical reasoning capabilities comparable to those of RL-based methods. The final model attains an average score of 48.2 across five mathematical reasoning benchmarks, which is similar to the performance of Qwen2.5-Math-7B-Instruct and other RL-based approaches with the some training data conditions.
@@
    Multi-round DPO requires significantly fewer computational resource can even be executed on a single GPU. Our sampling and training are conducted on four A800 GPUs using data parallelism. The full pipeline can be run on a single 80GB GPU or lower. In our setup with 4 A800 GPUs, each sampling round took 2–2.5 hours with vLLM (Kwon et al., 2023), training took 1 hour per round, and the full process completed in about 80 hours. On a single GPU, the full process would take around 3 days.
@@ 3.5.2 More Analysis
To further analyze these differences, we examine reasoning patterns in Figure 5. In Qwen2.5-Math-7B, self-checking and confirmation mechanisms are present but inconsistently applied, often leading to repeated questioning rather than definitive verification. In contrast, Qwen2.5-Math-7B-Instruct follows a more standardized CoT approach, ensuring a structured reasoning process. Our model, Qwen2.5-7B-DPO-VP, exhibits a distinct reasoning pattern—typically confirming the final answer after a single verification. This behavior likely results from the VP-based filtering process, which refines the model’s ability to select and reinforce correct reasoning paths. As a result, the model develops greater confidence in its final outputs, effectively reducing unnecessary re-evaluation while maintaining robustness in problem-solving.
@@ 4 Related Works
Self-improvement primarily occurs during the training phase, where the model enhances its own performance by evaluating and refining its outputs without relying on external human supervision (Xi et al., 2024). One notable technique that facilitates self-improvement is reflection (Zelikman et al., 2022), where LLMs analyze their reasoning processes to identify and correct mistakes. It is generally believed that LLMs are incapable of achieving self-improvement without any external information (Huang et al., 2024; Tyen et al., 2024). Consequently, an increasing number of methods leverage external supervision signals help self-improvement, for example, outcome labels for mathematical or programming problems (McAleese et al., 2024; Gao et al., 2024), process reward models (Zhang et al., 2024b), value function (Wang et al., 2024a) or LLM-as-a-judge (Li et al., 2024; Zhang et al., 2025) frameworks. The most closely related contemporary works to ours are Iterative DPO (Pang et al., 2024) and ThinkPO (Yang et al., 2025). The former achieves self-improvement by generating diverse data through multiple iterations of self-evaluation, while the latter focuses on selecting positive samples that encourage deeper reasoning. In contrast, our work conducts a thorough investigation into data selection for DPO and the iterative improvement framework of generator-verifier models.
@@
In early 2025, with the release of DeepSeek-R1 (Guo et al., 2025), efficient post-training paradigms gained significant attention (Trung et al., 2024). SimpleRL-Zero (Zeng et al., 2025) achieved remarkable performance through RL, training on just 8K challenging mathematical problems with rule-based rewards. Xiong et al. (2025) integrated reflective vocabulary into generative RM to enable selective RL-based correction. PURE (Cheng et al., 2025) revisited stepwise reward credit assignment in RL. Logic-RL (Xie et al., 2025) refined the penalty function for formatting errors, leading to improved logical reasoning capabilities. From the data utilization perspective, LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrated that fine-tuning with carefully curated long-chain reflective data can effectively amplify the latent reasoning capabilities. From the reasoning paradigm perspective, CFT (Wang et al., 2025) encouraged models to develop critical thinking during SFT. Our work focuses on achieving RL-based reasoning with significantly fewer computational resources. We demonstrate that, with a strong base model, coarse-grained selection is enough for performance enhancement, achieving competitive results with reduced computational overhead.
@@ Appendix B Details of Processbench Evaluation
Therefore, an alternative approach is to separately define the F1 score for positive and negative samples (Hinojosa Lee et al., 2024). The former is equivalent to the commonly used F1 score, while the latter is slightly modified based on the above formula. Then, a weighted F1 score is calculated based on the proportion of positive and negative samples. The weighted-F1 score better reflects the model performance.
We define positive samples as those labeled as incorrect. The formula for calculating the weighted F1 score is as follows:
@@ F.2 Iterative DPO Can Be Viewed as a High Off-Policy Degree Online RL
During the period of conducting this research, we observed that leveraging iterative DPO for model optimization has already become a viable low-cost alternative to RL fine-tuning for some enterprises and research institutions when extreme performance is not the primary goal. We are delighted by this coincidence, as it indirectly validates that we have caught up with the research frontier in the field of LLM reasoning. We also extend our gratitude to Hanning Zhang (2025), Liang Wen (2025), and all open-source contributors in the LLM community for their selfless contributions to the open-source ecosystem!

### v2

@@ 1 Introduction
    Iterative Improvement of Generator and RM can be Achieved via Multi-Round DPO: By relabeling the RM with online data, the mutual enhancement of the generator and RM can be realized through multiple rounds of DPO. In contrast, traditional RL frameworks typically rely on a fixed RM, making it difficult to update the RM in an online manner.
@@
    Multi-Round DPO Achieves RL-Level Performance with Lower Computational Cost: Our final model DPO-VP, based on Qwen2.5-Math-7B, achieves comparable average performance to SimpleRL-Zero, PURE-VR and DPO-R1-Zero across 5 challenging benchmarks (48.2 vs 48.8/47.7/47.0), while requiring only a single 80GB GPU for training.
@@ 3.5.1 Setup and Main Results
  6. (f)
    Qwen2.5-7B-DPO-R1-Zero (Zhang et al., 2025a): A concurrent work that applied iterative DPO without any extra SFT data.
To align with existing RL methods, following Zeng et al. (2025) and Cheng et al. (2025), we select 8K Level 3-5 questions from the MATH as self-improvement data, and additionally evaluate on Olympiad-level benchmarks, including AMC23, AIME24, and OlympiadBench He et al. (2024). Furthermore, we conduct additional comparisons on two out-of-domain datasets to assess generalization performance. In addition, we compare with Qwen2.5-7B-DPO-R1-Zero (Zhang et al., 2025a) that applies iterative DPO on 200K questions (more than 8K questions) from the NuminaMath dataset (Li et al., 2024a). Moreover, we employ annealed sampling to enhance response diversity. Specifically, the sampling temperature $t$ is set to 0.7 for the first three epochs, increased to 1.0 during epochs 4–5, and further raised to 1.2 in the final epoch. This annealing schedule is motivated by the observation that model performance nearly converges after three epochs, and increasing the temperature in later stages allows for fine-tuning on more diverse samples, leading to additional gains.
Table 2 presents the pass@1 accuracy of different models, while Table 3 further analyzes training data consumption and GPU usage.
@@
    Multi-round DPO can achieve mathematical reasoning capabilities comparable to those of RL-based methods. The final model attains an average score of 48.2 across five mathematical reasoning benchmarks, which is similar to the performance of Qwen2.5-Math-7B-Instruct and other RL-based approaches with the some training data conditions. Compared to Qwen2.5-7B-DPO-R1-Zero, our model still achieves a higher average score. We hypothesize that this may be attributed to the increased response diversity introduced by the annealed sampling strategy.
@@
    Multi-round DPO requires significantly fewer computational resource can even be executed on a single GPU. Our sampling and training are conducted on four A800 GPUs using data parallelism. The full pipeline can be run on a single 80GB GPU or lower. In our setup with 4 A800 GPUs, each sampling round took 2–2.5 hours with vLLM (Kwon et al., 2023), training took 1 hour per round, and the full process completed in about 1 day. On a single GPU, the full process would take around 3 days.
@@ 3.5.2 More Analysis
To further analyze these differences, we examine reasoning patterns in Appenedix Figure 7. In Qwen2.5-Math-7B, self-checking and confirmation mechanisms are present but inconsistently applied, often leading to repeated questioning rather than definitive verification. In contrast, Qwen2.5-Math-7B-Instruct follows a more standardized CoT approach, ensuring a structured reasoning process. Our model, Qwen2.5-7B-DPO-VP, exhibits a distinct reasoning pattern—typically confirming the final answer after a single verification. This behavior likely results from the VP-based filtering process, which refines the model’s ability to select and reinforce correct reasoning paths. As a result, the model develops greater confidence in its final outputs, effectively reducing unnecessary re-evaluation while maintaining robustness in problem-solving.
Why Iterative DPO Leads to Improvement
Based on the rigorous theoretical analysis in Xiong et al. (2024), intuitively, iterative DPO leads to improvement because it allows the model to continually refine its policy through updated feedback, thereby reducing reward misalignment over time. Additionally, the batch exploration strategy promotes diverse sampling, which enhances generalization and mitigates overfitting to early-stage preferences. The performance gains of iterative DPO on math tasks stem from the inherent self-reflection ability in the base model and the iterative nature of the near-online RL setting, as further discussed in Appendix F.
@@ 4 Related Works
Self-improvement primarily occurs during the training phase, where the model enhances its own performance by evaluating and refining its outputs without relying on external human supervision (Xi et al., 2024). It is generally believed that LLMs are incapable of achieving self-improvement without any external information (Huang et al., 2024; Tyen et al., 2024). Consequently, an increasing number of methods leverage external supervision signals help self-improvement, for example, outcome labels for mathematical or programming problems (McAleese et al., 2024; Gao et al., 2024), process reward models (Zhang et al., 2024b), value function (Wang et al., 2024a) or LLM-as-a-judge (Li et al., 2024b; Zhang et al., 2025b) frameworks. The most closely related contemporary works to ours are (Xiong et al., 2024), which proposed the iterative DPO with a rigorous theoretical analysis to demonstrate the advantage of online iterative training in the context of chat task. They further generalize the iterative DPO to the multi-turn scenario and take the potential external messages into consideration to tool-integrated reasoning tasks (Xiong et al., 2025a). In contrast, we conduct a preliminary study on data selection strategies in DPO and propose an iterative training framework for generator-verifier models, enabling mutual enhancement between the generator and the RM.
@@
In early 2025, with the release of DeepSeek-R1 (Guo et al., 2025), efficient post-training paradigms gained significant attention (Trung et al., 2024). SimpleRL-Zero (Zeng et al., 2025) achieved remarkable performance through RL, training on just 8K challenging mathematical problems with rule-based rewards. Xiong et al. (2025b) integrated reflective vocabulary into generative RM to enable selective RL-based correction. PURE (Cheng et al., 2025) revisited stepwise reward credit assignment in RL. Logic-RL (Xie et al., 2025) refined the penalty function for formatting errors, leading to improved logical reasoning capabilities. From the data utilization perspective, LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrated that fine-tuning with carefully curated long-chain reflective data can effectively amplify the latent reasoning capabilities. From the reasoning paradigm perspective, CFT (Wang et al., 2025) encouraged models to develop critical thinking during SFT. Our work focuses on achieving RL-based reasoning with significantly fewer computational resources. We demonstrate that, with a strong base model, coarse-grained selection is enough for performance enhancement, achieving competitive results with reduced computational overhead. Our overall training framework is more resource-efficient than the widely adopted SFT+RL pipeline, while outperforming concurrent iterative DPO methods.
@@ Appendix B Details of Processbench Evaluation
Therefore, an alternative approach is to separately define the F1 score for positive and negative samples (Hinojosa Lee et al., 2024). The former is equivalent to the commonly used F1 score, while the latter is slightly modified based on the above formula. Then, a weighted F1 score is calculated based on the proportion of positive and negative samples. The weighted-F1 score better reflects the model performance. We define positive samples as those labeled as incorrect. The formula for calculating the weighted F1 score is as follows:
@@ F.2 Iterative DPO Can Be Viewed as a High Off-Policy Degree Online RL
During the period of conducting this research, we observed that leveraging iterative DPO for model optimization has already become a viable low-cost alternative to RL fine-tuning for some enterprises and research institutions when extreme performance is not the primary goal. We are delighted by this coincidence, as it indirectly validates that we have caught up with the research frontier in the field of LLM reasoning. We also extend our gratitude to Zhang et al. (2025a), Liang Wen (2025), and all open-source contributors in the LLM community for their selfless contributions to the open-source ecosystem!